OpenStudy (anonymous):

Calculate the derivatives \(sin^{'}(x) = \frac{d}{dx}sin(x)\) and \(cos^{'}(x) = \frac{d}{dx}cos(x)\) by differentiating the summands of the power series. Show as a corollary the identity \(sin^{2}(x)+cos^{2}(x) = 1\) for all \(x \in \mathbb{R}\).

OpenStudy (anonymous):

i get such a help but i dont know how to apply it to question "the derivative of each of the terms then simplify and see what the power series now resembles "

OpenStudy (kinggeorge):

So the power series for \(\sin(x)\) is \[\large \sum_{n=0}^\infty\frac{(-1)^nx^{2n+1}}{(2n+1)!}=x-\frac{x^3}{3!}+\frac{x^5}{5!}-\frac{x^7}{7!}+...\]And the power series for \(\cos(x)\) is\[\large \sum_{n=0}^\infty\frac{(-1)^nx^{2n}}{(2n)!}=1-\frac{x^2}{2!}+\frac{x^4}{4!}-\frac{x^6}{6!}+...\]

OpenStudy (kinggeorge):

Now, look at \[x-\frac{x^3}{3!}+\frac{x^5}{5!}-\frac{x^7}{7!}+...\]what is the derivative of this?

OpenStudy (anonymous):

hmm what would you say what the derivate of this?

OpenStudy (kinggeorge):

Try to use the power rule on that to find the derivative. That is, \[\frac{d}{dx}(x^n)=nx^{n-1}\]Once you've looked at that, try to do it in general. What is the derivative of \[\frac{(-1)^nx^{2n+1}}{(2n+1)!}\]

OpenStudy (anonymous):

i am so bad i cant start with anthing George..

OpenStudy (anonymous):

\[(-1)^{n} \] is not convergenze i know :)

OpenStudy (kinggeorge):

It's not too bad. Since you're pressed for time, let's start with the derivative of \[\frac{(-1)^nx^{2n+1}}{(2n+1)!}\]Using the power rule (we can ignore the -1), we get that\[\frac{d}{dx}\left(\frac{(-1)^nx^{2n+1}}{(2n+1)!}\right)=\frac{(2n+1)(-1)^nx^{2n}}{(2n+1)!}\]You can cancel the \(2n+1\) from the top and bottom to get\[\frac{(-1)^nx^{2n}}{(2n)!}\]This is exactly the power sum for \(\cos(x)\), so this shows that \(\frac{d}{dx}(\sin(x))=\cos(x)\).

OpenStudy (anonymous):

ok..

OpenStudy (anonymous):

this end solution George ? or you want to add more things?

OpenStudy (anonymous):

i think it was end solution thank you very much @KingGeorge

OpenStudy (kinggeorge):

Now for the other part\[\frac{d}{dx}\left(\frac{(-1)^nx^{2n}}{(2n)!}\right)=\frac{2n(-1)^nx^{2n-1}}{(2n)!}\]Once again, we can cancel the \(2n\), and get \[\frac{(-1)^nx^{2n-1}}{(2n-1)!}.\]Now, notice what happens if we subsitute \(n\) for \(n+1\). We get \[-\frac{(-1)^nx^{2n+1}}{(2n+1)!}\]

OpenStudy (anonymous):

aah ok..

OpenStudy (kinggeorge):

The sum of this is exactly equal to \(-\sin(x)\), so we have proved that \(\frac{d}{dx}(\cos(x))=-\sin(x)\).

OpenStudy (anonymous):

ok thank you very much George, you are my live saver again.. i have 3 or 4 more questions i will post above, they are also not so difficult ones like the other weeks, if you like you can look at them

OpenStudy (kinggeorge):

I can take a look at them. Also, we still haven't done the corollary, but here's how to do it. Take the derivative of \(\sin^2(x)+\cos^2(x)\). Show that this is equal to 0. After that, substitute \(x=0\), to get a value of 1. Since the derivative is 0, the function is constant, and since it equals 1 at \(x=0\), it must be that \(\sin^2(x)+\cos^2(x)=1\).

OpenStudy (anonymous):

sorry i thought its finished, pls feel free to write more if we are still not finished with it

OpenStudy (kinggeorge):

The derivative of \(\sin^2(x)+\cos^2(x)\) is \[2\sin(x)\cos(x)+2\cos(x)(-\sin(x))=2\sin(x)\cos(x)-2\sin(x)\cos(x)=0\]

OpenStudy (anonymous):

ok..

OpenStudy (kinggeorge):

Then, at \(x=0\) we have \[\sin^2(0)+\cos^2(1)=0+1^2=1\]Thus, we have the identity \[\sin^2(x)+\cos^2(x)=1\]Since it must be constant for all \(x\in\mathbb{R}\)

OpenStudy (anonymous):

also sin value is always like zero

OpenStudy (kinggeorge):

The sin value is not always 0. However, the value of \(\sin^2(x)+\cos^2(x)\) is always 1. Remember, we were differentiating the entire function, and not just \(\sin(x)\).

OpenStudy (anonymous):

ok..

OpenStudy (kinggeorge):

Also, something just came up, and I'll have to be gone for about 20 minutes. So for now, you might want to get someone else to help you since I won't be here for a bit.

OpenStudy (anonymous):

and this is our end solution?

OpenStudy (kinggeorge):

Yes, this is the end solution.

OpenStudy (anonymous):

ok thank you very much George, i write them clear on paper now